# Unitary discriminants of characters

### v1

@@ 9 An example: The sporadic simple O’Nan group
$\begin{array}[]{|c|c|c|c|c|}\hline\cr\chi&\chi(1)&{\mathbb{Q}}(\chi)&\operatorname{disc}(\chi)&\Delta(\chi)\\ \hline\cr 3,4&13376&{\mathbb{Q}}[\sqrt{-31}]&1&[{\mathbb{Q}}]\\ \hline\cr 5,6&25916&{\mathbb{Q}}[\sqrt{-5}]&1&[{\mathbb{Q}}]\\ \hline\cr 31-34&342&{\mathbb{Q}}[\sqrt{2},\sqrt{-3}]&-1&[(-1,-1)_{{\mathbb{Q}}[\sqrt{2}]}]\\ \hline\cr 45-48&52668&{\mathbb{Q}}[\sqrt{7},\sqrt{-3}]&8+3\sqrt{7}&[(-3,8+3\sqrt{7})_{{\mathbb{Q}}[\sqrt{7}]}]\\ \hline\cr 53,54&63612&{\mathbb{Q}}[\sqrt{-3}]&55&[(-3,55)_{{\mathbb{Q}}}]\\ \hline\cr 57,58&116622&{\mathbb{Q}}[\sqrt{-3}]&-10&[(-3,-10)_{{\mathbb{Q}}}]\\ \hline\cr 59,60&122760&{\mathbb{Q}}[\sqrt{-3}]&1&[{\mathbb{Q}}]\\ \hline\cr 61-64&169290&{\mathbb{Q}}[\sqrt{2},\sqrt{-3}]&-1&[(-1,-1)_{{\mathbb{Q}}[\sqrt{2}]}]\\ \hline\cr 65,68&169632&{\mathbb{Q}}[\sqrt{15},\sqrt{-3}]&-2\sqrt{15}+15&[(-3,-2\sqrt{15}+15)_{{\mathbb{Q}}[\sqrt{15}]}]\\ 66,67&169632&{\mathbb{Q}}[\sqrt{15},\sqrt{-3}]&2\sqrt{15}+15&[(-3,2\sqrt{15}+15)_{{\mathbb{Q}}[\sqrt{15}]}]\\ \hline\cr 69,70&175770&{\mathbb{Q}}[\sqrt{-3}]&-11&[(-3,-11)_{{\mathbb{Q}}}]\\ \hline\cr 71,72&207360&{\mathbb{Q}}[c_{19},\sqrt{-3}]&3c_{19}^{2}+10c_{19}+9&[(-3,3c_{19}^{2}+10c_{19}+9)_{{\mathbb{Q}}[c_{19}]}]\\ 75,76&207360&{\mathbb{Q}}[c_{19},\sqrt{-3}]&-13c_{19}^{2}+3c_{19}+76&[(-3,-13c_{19}^{2}+3c_{19}+76)_{{\mathbb{Q}}[c_{19}]}]\\ 73,74&207360&{\mathbb{Q}}[c_{19},\sqrt{-3}]&10c_{19}^{2}-13c_{19}-29&[(-3,10c_{19}^{2}-13c_{19}-29)_{{\mathbb{Q}}[c_{19}]}]\\ \hline\cr 77-80&253440&{\mathbb{Q}}[\sqrt{93},\sqrt{-3}]&1&[{\mathbb{Q}}[\sqrt{93}]]\\ \hline\cr\end{array}$
@@
Also $\chi_{5},\chi_{6}=\overline{\chi_{5}}$ are non-faithful. Their character field is $L={\mathbb{Q}}[\sqrt{-5}]$, has class number 2, and is ramified at 2 and 5. The $p$-modular reduction of these two characters is unitary stable for all $p\neq 3,7$. As $3$ and $7$ are decomposed in $L/{\mathbb{Q}}$, Theorem 5.4 yields that the only primes that are possibly ramified in the discriminant algebra are $2$ and $5$. The orthogonal discriminant of the $5$-modular reduction of $\chi_{5}$ is a square mod 5, Corollary 5.6 implies that the prime 5 is not ramified in $\Delta(\chi_{5})$. As the number of ramified primes is even, the prime $2$ cannot be the only ramified prime, so $\Delta(\chi_{5})=[{\mathbb{Q}}]$.
@@
$\begin{array}[]{|l|c|c|c|c|c|c|c|c|c|c|}\hline\cr\chi&\dim&37&43&61&67&73&79&97&103&109\\ \hline\cr 45&72&yn&n&y&y&y&n&n&yn&yy\\ \hline\cr 57&162&yy&yy&nn&yy&nn&yy&nn&nn&yy\\ \hline\cr\end{array}$
@@
The character $\chi_{59}$ reduces absolutely irreducible modulo 5 and 11 leaving $1$ and $2$ as possible unitary discriminants. Its reduction modulo 3 is not orthogonally stable, so we cannot obtain any more information about this character from the decomposition matrix. So we use the unitary condensation method described in Section 8.2. We start with the permutation module $P$ of dimension $736560$ on the cosets of the subgroup $L_{3}(7)\leq 3.ON$ and use fixed point condensation with the 7-Sylow subgroup $H$ of $3.ON$. The group $H$ has 2196 orbits on the 736560 cosets, falling into 732 orbits under the center $\langle z\rangle$ of $3.ON$. Computing modulo primes $p\equiv 1\pmod{3}$ we specify the action of $z$ as some primitive third root of unity in ${\mathbb{F}}_{p}^{\times}$ and hence reduce the computations to find composition factors of a 732 dimensional ${\mathbb{F}}_{p}$-module for in total 25 suitable primes. These composition factors have dimension $3,171,192,366$, where the latter two correspond to characters number $53/54$ resp. $59/60$. For some $\alpha\in N_{3.ON.2}(H)$ of order $2$ that is not contained in $3.ON$ we compute the determinants of an $\alpha$-fixed skew-adjoint element in the condensed algebra as $3\cdot 5\cdot 11$ on the composition factor of dimension 192 and $7$ for the one of dimension 366. On the orthogonal complement of the fixed space the relevant discriminants of the restriction to the normaliser of $H$ are powers of $7$, so one concludes that $\operatorname{disc}(\chi_{59})=1$.
@@
Number 65-68 have character field $L:={\mathbb{Q}}[\sqrt{15},\sqrt{-3}]$ an unramified extension of its maximal real subfield $K={\mathbb{Q}}[\sqrt{15}]$. Both fields, $L$ and $K$ have class number 2. The fundamental unit of $K$ is a norm. The prime ideals of $K$ that are inert in $L/K$ and divide the group order are the divisors of $2$, $5$, $11$, and $19$. The reduction modulo both prime ideals of $K$ that divide 19 is absolutely irreducible. For 11, one prime ideal yields an absolutely irreducible reduction, the other one a non unitary stable reduction. By Proposition 5.7 exactly one of these two prime ideals, say $\wp_{11}$, ramifies in $\Delta(\chi_{65})$. Also the reduction modulo the unique prime ideal $\wp_{5}$ of $K$ that divides $5$ is not unitary stable. The ideal $\wp_{3}\wp_{5}\wp_{11}=(2\sqrt{15}+15)$ has a totally positive generator, but no such combination involving the prime ideal dividing 2 has such a totally positive generator. So we conclude that the discriminant of $\chi_{65}$ is represented by ±$\pm 2\sqrt{15}+15$.
@@
The characters number $71-76$ are algebraic conjugate and so are their unitary discriminants. It is enough to treat $\chi_{71}=\overline{\chi_{72}}$. The character field is $L=K[\sqrt{-3}]$ where $K={\mathbb{Q}}[c_{19}]$ is the real subfield of $L$. The extension $L/K$ is only ramified at 3, $L$ has class number 3, $K$ has class number 1 and all totally positive units in $K$ are squares. The prime divisors of the group order that are inert in $L/K$ are $2,5$ and the three prime ideals of $K$ that divide 11. Whereas $\chi_{71}\pmod{2}$ and $\pmod{5}$ is absolutely irreducible and hence unitary stable, its $11$-modular reduction is unitary stable at the prime ideal of $K$ that contains $c_{19}-6$ and not unitary stable at the other two prime ideals. So the unitary discriminant is a totally positive generator of the product of these two prime ideals dividing 11 in $K$. One checks that this number is indeed a square modulo 3, as predicted by the orthogonal discriminant of $\chi_{71}\pmod{3}$.

### v2

@@ 11 An example: The sporadic simple O’Nan group
[…]
@@
Also $\chi_{5},\chi_{6}=\overline{\chi_{5}}$ are non-faithful. Their character field is $L={\mathbb{Q}}[\sqrt{-5}]$, has class number 2, and is ramified at 2 and 5. The $p$-modular reduction of these two characters is unitary stable for all $p\neq 3,7$. As $3$ and $7$ are decomposed in $L/{\mathbb{Q}}$, Theorem 6.3 yields that the only primes that are possibly ramified in the discriminant algebra are $2$ and $5$. The orthogonal discriminant of the $5$-modular reduction of $\chi_{5}$ is a square mod 5, Corollary 6.5 implies that the prime 5 is not ramified in $\Delta(\chi_{5})$. As the number of ramified primes is even, the prime 22 cannot be the only ramified prime, so $\Delta(\chi_{5})=[{\mathbb{Q}}]$.
@@
The characters number 313431-34 restrict irreducibly to the maximal subgroup $C_{3}\times L_{3}(7).2$. Their unitary discriminant has been computed by David Schlang by restricting further to the normaliser of a 7-Sylow subgroup of $L_{3}(7)$.
@@
The character $\chi_{59}$ reduces irreducibly modulo 5 and 11 leaving $1$ and $2$ as possible unitary discriminants. Its reduction modulo 3 is not orthogonally stable, so we cannot obtain any more information about this character from the decomposition matrix. So we use the unitary condensation method described in Section 10.2. We start with the permutation module $P$ of dimension $736560$ on the cosets of the subgroup $L_{3}(7)\leq 3.ON$ and use fixed point condensation with the 7-Sylow subgroup $H$ of $3.ON$. The group $H$ has 2196 orbits on the 736560 cosets, falling into 732 orbits under the center $\langle z\rangle$ of $3.ON$. Computing modulo primes $p\equiv 1\pmod{3}$ we specify the action of $z$ as some primitive third root of unity in ${\mathbb{F}}_{p}^{\times}$ and hence reduce the computations to find composition factors of a 732 dimensional ${\mathbb{F}}_{p}$-module for in total 25 suitable primes. These composition factors have dimension $3,171,192,366$, where the latter two correspond to characters number $53/54$ resp. $59/60$. For some $\alpha\in N_{3.ON.2}(H)$ of order $2$ that is not contained in $3.ON$ we compute the determinants of an $\alpha$-fixed skew-adjoint element in the condensed algebra as $3\cdot 5\cdot 11$ on the composition factor of dimension 192 and $7$ for the one of dimension 366. On the orthogonal complement of the fixed space the relevant discriminants of the restriction to the normaliser of $H$ are powers of $7$, so one concludes that $\operatorname{disc}(\chi_{59})=1$.
@@
Number 65-68 have character field $L:={\mathbb{Q}}[\sqrt{15},\sqrt{-3}]$ an unramified extension of its maximal real subfield $K={\mathbb{Q}}[\sqrt{15}]$. Both fields, $L$ and $K$ have class number 2. The fundamental unit of $K$ is a norm. The prime ideals of $K$ that are inert in $L/K$ and divide the group order are the divisors of $2$, 55, $11$, and $19$. The reduction modulo both prime ideals of $K$ that divide 19 is irreducible. For 11, one prime ideal yields an irreducible reduction, the other one a non unitary stable reduction. By Proposition 6.6 exactly one of these two prime ideals, say $\wp_{11}$, ramifies in $\Delta(\chi_{65})$. Also the reduction modulo the unique prime ideal $\wp_{5}$ of $K$ that divides $5$ is not unitary stable. The ideal $\wp_{3}\wp_{5}\wp_{11}=(2\sqrt{15}+15)$ has a totally positive generator, but no such combination involving the prime ideal dividing 2 has such a totally positive generator. So we conclude that the discriminant of $\chi_{65}$ is represented by $\pm 2\sqrt{15}+15$.
@@
The characters number 717671-76 are algebraic conjugate and so are their unitary discriminants. It is enough to treat $\chi_{71}=\overline{\chi_{72}}$. The character field is $L=K[\sqrt{-3}]$ where $K={\mathbb{Q}}[c_{19}]$ is the real subfield of $L$. The extension $L/K$ is only ramified at 3, $L$ has class number 3, $K$ has class number 1 and all totally positive units in $K$ are squares. The prime divisors of the group order that are inert in $L/K$ are $2,5$ and the three prime ideals of $K$ that divide 11. Whereas $\chi_{71}\pmod{2}$ and $\pmod{5}$ is absolutely irreducible and hence unitary stable, its $11$-modular reduction is unitary stable at the prime ideal of $K$ that contains $c_{19}-6$ and not unitary stable at the other two prime ideals. So the unitary discriminant is a totally positive generator of the product of these two prime ideals dividing 11 in $K$. One checks that this number is indeed a square modulo 3, as predicted by the orthogonal discriminant of $\chi_{71}\pmod{3}$.